# Robust surface superconductivity and vortex lattice in the Weyl semimetal -PtBi2

### v1

@@ IV DFT calculations
V nanometer-sized flakes
We have imaged the surface of our cleaved samples after STM measurements using high-resolution Scanning Electron Microscopy (SEM) to characterize the origin and composition of the nanometer-sized flakes. In Fig. 5 (a) we show a typical image of several microns showing several step edges and 120∘ corners, reflecting the trigonal symmetry of the crystal lattice. In Fig. 5 (b) we show a zoom on a flat terrace showing flakes of the length scale we have observed in our STM measurements. We have performed Energy Dispersive electron Spectroscopy (EDS) measurements on some of these objects and on top of the flat defect-free surface to analyze the elemental composition. EDS essentially yields the same composition for all measured spots, i.e, Pt:Bi= which coincides with the stoichiometric ratio within the experimental uncertainty of . Many of these flakes are found around step edges and some have a rough hexagonal shape. Possibly these are caused by defects during cleavage.
VI Vortices in PtBi2
VI.1 Magnetic field history
The observed behavior (absence of a clear vortex lattice and instead the presence of a homogeneous tunneling conductance on atomically flat terraces and vortex lattice in nanometer-sized flakes) is independent on the magnetic field history (field cooling, zero field cooling), the reversal of the direction of the magnetic field, the reversal of the setup bias voltage and the size of the tunneling current down to 0.1 nA.
VI.2 Interaction with nanometer-sized flakes
In Figs. 6,7,8 we present three experiments which we consider interesting, because they provide more information about the nature of the vortex lattice in the nanometer-sized flakes.
In Fig. 6(a) we show a zero-bias conductance map of a single vortex together with the simultaneously-acquired topography in Fig. 6(c). We observe that the vortex is confined to the nanometer-sized flake. In Fig. 6(b,d) we plot the tunneling conductance and topography profiles as a function of the position as the tip moves from the nanometer-sized flake (delimited with white dashed lines in Fig. 6(a,c) onto the atomically flat regions). We see in Fig. 6(b) a fully opened superconducting gap on the atomically flat terrace and a neat vortex on the nanometer-sized flake.
Another example of an atomically flat surface at an applied magnetic field of 0.5 T is shown in Fig. 7(a). This surface presents a hole on the upper left corner. In Fig. 7(b) we see the result of a first scan. Inside the hole we find a nanometer-sized flake. Although the superconducting gap is perfectly homogeneous over the whole surface, we see the vortices on the nanometer-sized flake within the hole Fig. 7(c). After making several scans, we find the topography shown in Fig. 7(e). The surface inside the hole is now perfectly flat, because the tip motion has changed the position of the nanometer-sized flake, possibly removing it altogether from this field of view. We then see no trace of the vortices Fig. 7(f) and instead a spatially homogeneous superconducting tunneling conductance with the properties discussed in the main text.
Finally, in Fig. 8 we elaborate further on vortex dynamics on the field of view shown in Fig. 3 (e,f) of the main text. In Fig. 8 (a) we show the topography acquired simultaneously with the tunneling conductance map shown in Fig. 3(e) of the main text. In the center of the field of view there is a large nanometer-sized flake. On the bottom left corner (orange square) there is a stripe-like flake separating two atomically flat regions. In Fig. 8(b) we show the tunneling conductance map corresponding to this area. We see that vortices appear blurred and confined to a central line. It is important to note that this striple-like nanometer-sized flake is smaller in height than the one present on the center of Fig. 8(a). The height profile, shown in Fig. 8(c,d) shows a difference in size of more than an order of magnitude. This implies that vortices length is smaller on the stripe-like nanometer-sized flake, which allows for vortex mobility and produces a one-dimensional blurred feature in the tunneling conductance map. This suggests that there is a gradual evolution with increasing flake height from the free-moving vortices on flat surfaces to the completely pinned vortices on big nanometer-sized flakes.
It remains unclear why the effective length of superconducting vortices extends over larger length scales along the c-axis on flakes as compared to atomically flat areas. PtBi2 consists of weakly coupled layers of PtBi2 groups, separated by a van der Waals gap [18]. It could be that the areas where we observe vortices are composed of disordered groups of weakly coupled flakes. Each group will contain surface superconductivity and carry vortices, which interact through their magnetic field. This might allow for the establishment of interactions among two-dimensional surface vortices. We remind that, similarly, pancake vortices are observed with STM on Bi2Sr2CaCu2Ox. There are no reports of electrostatic STM tip-vortex interactions with two-dimensional vortices in Bi2Sr2CaCu2Ox. The magnetic interaction among pancakes at different layers allows for vortex stability and their observation with STM [80, 81].
VI.3 Vortex dynamics temperature dependence
We show the effect of increasing temperature on the vortex lattice in Fig. 9(a-d). The scans are made consecutively and in the same field of view. The nanometer-sized flake slightly changes is lateral size with subsequent scans, but the surface remains large enough to identify the vortex lattice at the applied magnetic field (0.7 T). We see a hexagonal vortex lattice at 0.1 K, which remains but is considerably disordered at 1.2 K. At 1.5 K, there are practically no traces of superconducting vortices, although the average tunneling conductance normalized to its value at high voltages is well below one. This shows that vortices are moving below the tip. At 2 K, the tunneling conductance on the nanometer-sized flake is homogeneous, suggesting thermally activated vortex motion.
Remarkably, the tunneling conductance outside the nanometer-sized flake remains below its value inside. This suggests that vortices have a different dynamics on atomically flat surfaces and on the nanometer-sized flake.
VI.4 Vortex core size as a function of the magnetic field
The extension of vortices is related to the order parameter spatial dependence and, through this, to the coherence length . A natural model to gauge the vortex core size is thus to consider the order parameter slope at the vortex core center, such that . This model is described in Ref. [82]: In a STM measurement we can relate the spatial distribution of the order parameter with the normal DOS and the tunneling conductance using deGennes model [83] and considering that the order parameter must evolve smoothly, i.e., at half of the intervortex distance L due to the periodicity of the vortex lattice. The resulting expression is:
where , , and is the radius of the circular cell we use to approximate the hexagonal Wigner-Seitz cell of the vortex lattice [82]. In Fig. 10 we show the vortex core size , as obtained from fitting the radially-averaged tunneling conductance along vortex cores to equ. 1 following Ref. [82]. We find that the vortex core size is slightly decreasing with the magnetic field and tends towards the coherence length nm at the critical field T. This behavior is expected for a superconductor with an electronic mean free path of order of [82], confirming the strong reduction of the mean free path on nanometer-sized flakes.
VII Electrostatic interactions between tip and vortex lattice
We illustrate the interaction between tip and sample in Fig. 11. As is well known, a STM is most often blunt, but there is always a small atomic size tip which dominates in the tunneling process, because of the exponential dependence of the tunneling current with distance [84]. We represent the tunneling tip by a small triangular-like tip in Fig. 11. The tunneling current from this tip is shown schematically in red. The overall shape of the tip is however blunt and we represent it by a semi-spherical object with a radius of several nm. Indeed, the electrostatic interaction between tip and sample occurs over a much larger scale than tunneling, which depends on the size of the tip but is typically several nm in diameter [85, 86, 87, 88, 89, 90]. The electrostatic interaction between tip and dipoles on the sample contains several components, such as image charges and the actual dipolar interaction [85, 89, 90]. The latter can be, in principle, attractive and repulsive, depending for instance on the sign of the bias. Independently on the sign of the interaction, for a certain electrostatic configuration, there is an equilibrium distribution of vortices below the tip, most often consisting of the vortex lattice confined within the geometry of the capacitor built by the spherical tip (with nm or tens of nm radius). This equilibrium distribution of vortices below the tip is carried when moving the tip.
In practice, the interaction can just lead to changes in the position of a few vortices, because the hexagonal vortex lattice is not perfect and the vortex positions are slightly influenced by small defects. With very weak pinning, we obtain the result shown schematically in Fig. 11(b,c). When we move the tip the equilibrium position some vortices move together with the tip in such a way as to maintain the distribution of charges below the tip. The crux is that the point at which we tunnel is rarely, coinciding with a vortex core, as it is not located at the center of the capacitor built by the tip and the sample. Therefore, in most cases, no signature of vortices can be found from tunneling at the tip’s apex.
By contrast, adding additional two-dimensional superconducting layers on top of nanometer-sized flakes, as schematically represented in Fig. 11(d), enhances the effective length of the magnetic vortex. This increases the pinning force, eventually leading to . In that case, we have the situation proposed in Fig. 11(e,f). The vortex lattice on the flakes remains fixed (dark blue on Fig. 11(e,f), ), but vortices on the flat two-dimensional surface are moved by the interaction with the tip (light blue on Fig. 11(e,f), ).
In a few cases, we observe indeed changes in the density of states on the flat regions. Furthermore, at elevated temperatures and under magnetic fields (Fig. 9(d)), the zero bias tunneling conductance on the atomically flat areas at 0.7 T is higher than the one observed at zero magnetic field and at the temperature of the measurement ( and , where is the tunneling conductance normalized to its value at high bias). This shows that the atomically flat areas are interspersed with fluctuating vortices with increased temperature.
VIII STM tunneling with a superconducting tip
To obtain further evidence of the robustness of the superconducting properties of -PtBi2 we performed controlled indentations of the tip into the sample, picking up pieces of PtBi2. We observed the Josephson effect and traced its evolution when changing the tunneling current. We observed similar behavior as found previously in systems as Pb or Al [64]. There are peaks in the tunneling conductance at twice the superconducting gap and at the superconducting gap, due to Andreev reflection. In addition, there is a pronounced peak at zero bias which evidences the passage of a Cooper pair (Josephson) supercurrent (Fig. 12 (a)). When measuring with a tip of Pb, we also observe a Josephson peak (Fig. 12 (b)) and the features in the tunneling conductance characteristic of tunneling between superconductors with different gap sizes [94].

### v2

@@ VIII DFT calculations
The cleaving energies calculated within PBE+D3 [107, 119] for each of the planes of the crystal structure along the c-axis, shown in Fig. 1(a) (middle panel) are of 2.66 J/m2 for Pt-Bi2, 3.01 J/m2 for Pt-Bi3 and 0.89 for Bi1-Bi3. This agrees with the structural parameters, which provide an interlayer Pt-Bi3 distances of 1.33 Å, but an interlayer Bi1-Bi3 distance of 2.36 Å. Therefore, the cleaving plane is in between the Bi1 and Bi3 layers, giving the two surfaces shown in Fig. 1(b,c) of the main text.
IX Comparison with the superconducting properties found in literature
Previous bulk resistivity measurements have found a transition to the superconducting state at 0.6 K-1.1 K[120, 14]. However, the critical current and the critical field were very small. No trace of superconductivity is found in the specific heat down to 0.5 K [16]. Thin flakes with a thickness of the order of tens of nanometers have shown a Berezinskii-Kosterlitz-Thouless transition at mK [35]. Superconductivity with up to K appears with doping with Rh, Se and Te[120, 121] or by applying hydrostatic pressure [36].
On the other hand, techniques probing the surface have found a plethora of different critical temperatures ranging from K (point-contact) [15], to K (SQUID on tip) [16] and to 8-14 K [13] (ARPES). Most remarkably, STM measurements found superconducting gaps compatible with a of the order of 80 K[16, 37] and a T. However, superconducting properties probed in Ref. [16] were extremely inhomogeneous and even contained non-superconducting regions [18]. Furthermore, the vortex lattice could not be observed in Ref. [16] despite having performed measurements under high magnetic fields.
Other work in -PtBi2 did not find evidence for superconductivity in STM [17] and ARPES (down to 3 K) [33]. This is compatible with the very recent experiment of Ref.[52], where bulk superconductivity with a very small upper critical field and vortices whose size is an order of magnitude larger than the ones we report here, were observed below 0.35 K. Furthermore, a surface superconducting density of states below 2.9 K was observed in atomically flat areas similar to those we report in our experiment, again without observing the two-dimensional surface vortex lattice.